# Interpretable Modeling of Articulatory Temporal Dynamics from real-time MRI for Phoneme Recognition

### v1

@@ 2.2 Spatial-temporal dynamics via Optical Flow
Optical flow [16] is an alternative representation for modeling temporal dynamics in video sequences. Traditional optical flow methods [17] capture pixel-level motion but produce noisy representations at the object level, particularly in rtMRI video. While no deep learning models have been trained on medical images due to the lack of ground-truth annotations, our qualitative evaluation suggests that MemFlow [16], whose checkpoint was pretrained on the Sintel dataset, captures object-level motion patterns in rtMRI to some extent. Therefore, we extract optical flow fields $\mathcal{F}_{t}(x)\in\mathbb{R}^{T,C,W,H}$ from video set $v$, where C = 2 represents horizontal and vertical motion components. To stabilize noise in the estimated optical flow, we suppress irrelevant background motion by applying a mask derived from intensity thresholding of the original images, thereby retaining only articulator-related movements. The masked flows are subsequently encoded with a Vision Transformer [18] to obtain spatial flow embeddings $e_{\text{flows}}\in\mathbb{R}^{T,D}$. Temporal dependencies are then modeled using a temporal encoder (Mamba/LSTM), and the output embeddings are passed through a lightweight classification head to generate phoneme probability distributions in a way identical to the ROI approach (Figure  1).
@@
Raw Video ViT-B ❅
LSTM +
Conformer
0.49 N/A
@@ 4.1 Comparison of features
PER and per-phoneme classification accuracy (Top-1 and Top-3) for each speech representation are shown in Table  1. Since predicted phoneme sequences with the same rtMRI corpus [6], we include their results for comparison although our testing dataset is different. Models trained independently on raw video, optical flow, and six-channel ROI have the best PER values of 0.37, 0.41, and 0.53 respectively. While Top-1 per-phoneme classification accuracy remains consistent for all features, Top-3 per-phoneme classification accuracy is best with raw video and optical flow models.
The results for PER and per-phoneme classification accuracy of multimodal models trained on pairs of rtMRI-acquired features are also presented in Table  1. We also present confusion matrices for our best-performing model trained on six-channel ROI and raw video in Figure  3. These matrices show that most phonemes are recognized to varying degrees apart from /Z/ and /S/. While Top-3 per-phoneme classification accuracy fails to exceed that of single feature models, Top-1 per-phoneme classification accuracy is improved by using multiple features. The lowest PER of 0.34 was acquired with a multimodal model trained on six-channel ROI and raw video, indicating that these two feature types provide complementary information.

### v2

@@ 2.2 Spatial-temporal dynamics via Optical Flow
Optical flow [5] is an alternative representation for modeling temporal dynamics in video sequences. Traditional optical flow methods [19] capture pixel-level motion but produce noisy representations at the object level, particularly in rtMRI video. While no deep learning models have been trained on medical images due to the lack of ground-truth annotations, our qualitative evaluation suggests that MemFlow [5], whose checkpoint was pretrained on the Sintel dataset [3], captures object-level motion patterns in rtMRI to some extent. Therefore, we extract optical flow fields $\mathcal{F}_{t}(x)\in\mathbb{R}^{T,C,W,H}$ from video set $v$, where C = 2 represents horizontal and vertical motion components. To stabilize noise in the estimated optical flow, we suppress irrelevant background motion by applying a mask derived from intensity thresholding of the original images, thereby retaining only articulator-related movements. The masked flows are subsequently encoded with a Vision Transformer [26] to obtain spatial flow embeddings $e_{\text{flows}}\in\mathbb{R}^{T,D}$. Temporal dependencies are then modeled using a temporal encoder (Mamba/LSTM), and the output embeddings are passed through a lightweight classification head to generate phoneme probability distributions in a way identical to the ROI approach (Figure  1).
@@
Raw Video ViT-B ❅ LSTM + Conformer 0.49 N/A
@@ 4.1 Comparison of features
PER and per-phoneme classification accuracy (Top-1 and Top-3) for each speech representation are shown in Table  1. Since [14] predicted phoneme sequences with the same rtMRI corpus, we include their results for comparison although our testing dataset is different. Models trained independently on raw video, optical flow, and six-channel ROI have the best PER values of 0.37, 0.41, and 0.53 respectively. While Top-1 per-phoneme classification accuracy remains consistent for all features, Top-3 per-phoneme classification accuracy is best with raw video and optical flow models.
The results for PER and per-phoneme classification accuracy of multimodal models trained on pairs of rtMRI-acquired features are also presented in Table  1. We also present confusion matrices for our best-performing model trained on six-channel ROI and raw video in Figure  3. These matrices show that most phonemes are recognized to varying degrees apart from /Z/ and /S/. The matrices also show errors between similar consonants (e.g., velars /k/ and /g/). While Top-3 per-phoneme classification accuracy fails to exceed that of single feature models, Top-1 per-phoneme classification accuracy is improved by using multiple features. The lowest PER of 0.34 was acquired with a multimodal model trained on six-channel ROI and raw video, indicating that these two feature types provide complementary information.
@@ 5 Conclusion and Future Work
6 Acknowledgement
This work was supported by The U.S. National Science Foundation under Grant NSF (IIS-2311676, BCS-2240349, RI-2106930). We would like to thank Prof. Sudarsana Reddy Kadiri for his valuable comments and insights.